OpenStudy (cwtan):

Help.... I've done using many ways but I still fail to prove that: \[\frac{tan(x)+sec(x)-1}{tan(x)-sec(x)+1}=tan(x)+sec(x)\] . Please help me............

OpenStudy (shubhamsrg):

multiply divide LHS by (tan x + sec x).. leave one in numerator..simply with denominator.. you should get your ans..

myininaya (myininaya):

did you do what @shubhamsrg did? Or you could do which is much longer: I write both sides in terms of cosine and sine Did you try that?

OpenStudy (campbell_st):

use fractions tan = sin/cos sec = 1/cos 1 = cos/cos so the numerator is \[\frac{\sin(x)}{\cos(x)} + \frac{1}{\cos(x)} - \frac{\cos(x)}{\cos(x)}\] this can be simplified to \[\frac{\sin(x) - \cos(x) + 1}{\cos(x)}\] same approach for the denominator then when dividing by a fraction take the reciprocal and multiply

OpenStudy (cwtan):

LHS multiply by\[\frac{tan(x)+sec(x)-1}{tan(x)-sec(x)+1}\times\frac{tan(x)+sec(x)}{tan(x)+sec(x)}\] like this?

OpenStudy (anonymous):

(tanx +secx -(sec^2x-tan^2x))/(tanx-secx+1) =(tanx+secx)(1-secx+tanx)/((tanx-secx+1) =(tanx+secx) ..

OpenStudy (shubhamsrg):

yes.. @matricked has done it perfectly right..

OpenStudy (cwtan):

Oh I get it... \[\frac{tan(x)+sec(x)-1}{tan(x)-sec(x)+1}\times \frac{tan(x)+sec(x)}{tan(x)+sec(x)}=\\ \frac{(tan(x)+sec(x)-1)(tan(x)+sec(x))}{tan^2(x)-sec^2(x)+tan(x)+sec(x)}=\\ \frac{(tan(x)+sec(x)-1)(tan(x)+sec(x)}{tan(x)+sec(x)-1}=tan(x)+sec(x)\]

OpenStudy (shubhamsrg):

bingo!!